The load on an installation is 800 kW. 0.8 lagging PF which works for 3000 hours per annum. The tariff is 1 $ per kVA plus 0.2 Cents per kWh. If the power factor is improved to 0.9 lagging by means of loss-free capacitors costing 0.6 $ per kVAR, calculated the annual saving effected. Allow 10% per annum for interest and depreciation on capacitors.


1
Expert's answer
2021-05-19T04:37:02-0400

Leading kVAR taken by the capacitor "P(tan \phi_1 -tan \phi_2)=800(0.75 -0.4843)=212.56"

Annual cost before P.F correction

kVA demand ="\frac{800}{0.8}=1000"

kVA demand charges "=\$1 \times 1000= \$1000"

Units consumed 1 year "= 800 \times 3000 =2400000kWh"

Energy charges 1 year ="0.002 \times 2400000=\$4800"

Total annual cost "= \$ (1000+4800)=\$5800"

Annual cost after correction.

kVA demand ="\frac{800}{0.9}=888.89"

kVA demand charges "=\$1 \times 1000= \$888.89"

Energy charges = same as before i.e $4800

Capital cost of capacitors ="\$0.6 \times212.56=\$127.536"

Annual interest and depreciation ="\$0.1 \times 127.536=\$12.7536"

Total annual cost ="\$(888.89+4800+12.7536)\$5701.6436"

Annual saving ="5800-5701.6436=\$98.3564"